How do you solve for the temperature of air using the following equation?

v=345m/s

(331)N 273K v= (331 T S
Answer #1

given v = 345 m/s

v = 331 m/s \sqrt{\frac{T}{273}}

put the value of v in equation

(\frac{345}{331}) = \sqrt{\frac{T}{273}}

squre on both sides

(\frac{345}{331})^2 = {\frac{T}{273}}

Solve for T

1.08638110276*273 = T

T = 296.6 K = temperature of air